Circular No. 7250

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999dq IN NGC 976 S. Jha, P. Garnavich, P. Challis and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 1999dq, taken by P. Berlind on Sept. 4.5 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a peculiar type-Ia supernova at a very early epoch. The spectrum exhibits a blue continuum and a shallow Si II (rest 635.5 nm) absorption, as well as strong Fe III features observed at 500 nm and 431 nm. The spectrum is similar to those of SN 1998es (IAUC 7054) and SN 1991T (Filippenko et al. 1992, Ap.J. 384, 15) and implies that SN 1999dq will have a slow declining lightcurve. Narrow Hα emission superimposed along the line of sight to the supernova yields a recession velocity of 4160 km/s for the host galaxy. The supernova photospheric expansion velocity derived from the Ca II H+K absorption is 20 400 km/s, implying that the supernova is ∼ 10 days before maximum light. The Galactic visual extinction toward NGC 976 based on Schlegel, Finkbinder and Davis (1998, Ap.J. 500, 525) maps is 0.34 mag. The spectrum also shows interstellar Na I D absorption from the Galaxy and NGC 976, with equivalent widths of 0.15 and 0.08 nm, respectively, indicating that the supernova is likely reddened. The spectrum of SN 1999dq is available at http://cfa-

www.harvard.edu/cfa/oir/Research/supernova/RecentSN.html.

SUPERNOVA 1999dp IN UGC 3046 Jha, Garnavich, Challis and Kirshner also report that a noisy spec-

trum of SN 1999dp, again obtained by Berlind as above, indicates that this supernova is of type II. The spectrum shows a flat continuum with broad P-Cyg profiles of Hα and He I, as well as hints of Hβ absorption. Superimposed narrow Hα emission implies a recession velocity of 4670 km/s for UGC 3046. The photospheric expansion velocity measured for the supernova from the Hα profile is 9700 km/s.

COMET C/1999 H1 (LEE) Total visual magnitude estimates: Aug. 6.08 UT, 6.6 (K. Hornoch,

Lelekovice, Czech Republic, 10 × 80 binoculars); 10.07, 7.8 (K. Sárneczky, Szatymaz, Hungary, 20 × 60 binoculars); 14.01, 6.7 (M. Reszelski, Szamotuly, Poland, 20 × 60 binoculars); 19.10, 8.0 (A. Kammerer, Ettlingen, Germany, 0.20-m reflector; 9′ antitail in p.a. 70^{o}); 22.05, 7.5 (Hornoch; 20′ antitail in p.a. 110^{o}); 25.11, 7.2 (Kammerer); 28.95, 7.5 (Reszelski); 30.78, 7.7 (A. Baransky, Kiev, Ukraine, 12 × 80 binoculars); Sept. 1.92, 7.6 (Reszelski); 4.97, 7.2 (J. Carvajal, Avila, Spain, 10 × 70 binoculars).

1999 September 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden